# SPrimary Studies

### v1

@@ 5 Findings
  • addressing privacy considerations entailed from using third-party apps or libraries [S28, S47],
  • evaluating the consistency between privacy policies and actual app code [S37, S52, S54, S60],
  • detecting abuse of data collection due to granted permissions [S42, S43, S49, S50]
We noticed that technical contributions were often built on existing tools. The novelty of these contributions resided in the integration of existing tools and processes into a comprehensive pipeline for answering certain research questions.
@@
The aforementioned technical contributions were in all cases accompanied with validation elements such as, e.g., proof of concept examples [S6], user studies [S16, S59], or empirical validation using datasets. These datasets vary in size from small with less than 50 elements such as apps, privacy policies, app reviews, or programming methods [S42]. Medium-sized datasets range from 51–500 elements [S16, S28, S30, S50, S52, S53, S57], while large datasets contain more than 500 elements [S1, S8, S20, S35, S36, S37, S41, S44, S46, S49, S51, S53, S54].
As part of this RQ, we also looked into the app stores and operating systems targeted in the literature. Our results show that 31 papers considered Android only, eight considered both Android and iOS, one paper considered Android and Windows systems, and one paper focused exclusively on iOS.
@@
By examining the 25 studies with technical contributions, we observe that all but four are exclusively focused on Android. Regarding the outliers, one utilizes Amazon Alexa (a virtual assistant) [S37], one is based on iOS [S59], and the last two mention Apple App Store as well as Google Play Store for work based on explicit consent (through GUI testing) and privacy policies [S49, S57]. The remaining 21 studies focusing on Android include: One study that investigates Xiaomi market Store [S43], five lack comprehensive details about app stores [S41, S42, S44, S50, S60], and the remaining ones utilize GooglePlay [S8, S16, S20, S28, S30, S35, S36, S41, S44, S46, S47, S51, S52, S53, S54].
@@
Out of the 60 papers, only 14 papers ($\approx$23%) have shared some artifacts associated with the conducted research; Table 6 lists these papers and indicates the type of artifacts (source code, datasets, output files) made available. Out of these 14 papers, only two papers released all three types of artifacts.
@@
Source code Datasets Output files Papers
[…]
@@ 5.5 Research Gaps
1. More research investigating personal data collection is needed. Table 4 shows that the primary studies have extensively investigated the collection of personal data and recipients of personal data sharing activities. These topics are often related to analyzing security issues concerning data leaks. To this end, we identify the following research gaps:
  • 1.1
@@
  • 1.2
    Nowadays, personal data collection typically involves permission requests to access personal data collected by other apps on the same mobile device. Another research gap concerns the investigation of indirect collection of personal data. It is important to understand what personal data is collected indirectly, how the process is documented, and whether users are aware of that.
2. Exploring the legal bases other than “consent’ can have an impact on mobile apps development. Our results show that most primary studies investigate the explicit consent of a user for data collection and processing. While consent is indeed an important legal basis, other legal bases are as important. For instance, collecting and processing personal data based on legitimate interest for the purpose of investigating a security attack is a good example where consent might not even be asked. The legal basis for handling personal data provides a complex, yet unexplored research landscape.
3. The elicitation of comprehensive implementation details of data subject rights is paramount for developing legally-compliant mobile apps. Our SLR highlights that the existing work investigated data subject rights and their implementation in mobile apps only to a very limited extent. According to the GDPR, users (as data subjects) should be able to easily exercise their rights to access their personal data, lodge a complaint, etc. For instance, mobile apps should provide adequate and explicit information (ideally through direct interactions) for users to request account deletion or to access all their personal data collected during a specific time span [44]. In-depth understanding of data subject rights should first be established in research to then make its way into practice. Without such an awareness level on the different users’ rights, developing better and legally-compliant apps is subject to the random, potentially correct interpretations of developers.
@@ 6 Threats to Validity
Internal Validity. The main consideration concerning the internal validity is selection bias. To mitigate the effect of this threat, we followed the guidelines reported in the literature for conducting systematic literature reviews. The results of our data extraction are also made publicly available and are thus open to scrutiny.
@@
External Validity. We have explicitly defined the time range of interest for our SLR on the GDPR. We also iteratively refined our search query and retrieved the primary studies from multiple online repositories and search engines. While we are unable to cross out the potential threat of missing primary studies, we are confident that our findings are representative of the SE engineering research landscape and can thus be generalized.
@@ 7 Conclusion
In this paper we have presented our systematic literature review (SLR) concerning the research landscape in software engineering (SE) on addressing GDPR-relevant privacy concerns in mobile apps. We have comprehensively reviewed 60 primary studies, and analyzed them with respect to an existing comprehensive conceptual model which characterizes all possible privacy-relevant information types according to the GDPR provisions
We found that existing work focuses on three main GDPR concerns, related to: (i) sharing personal data with other entities (e.g., third parties) outside mobile apps, (ii) obtaining explicit consent from users, and (iii) collecting from the users directly certain categories of personal data. Additionally, our results show that the majority of existing work investigates problems such as data leaks, app permission requests, and obtaining users’ consent. To address these challenges, existing work relies to a large extent on natural language processing, machine learning, and program analysis.
We further identified research gaps that are not yet well explored in the literature. These gaps include the necessity of: (i) consistency checking of what the apps disclose to users through privacy policies against what the apps’ implementation details, (ii) investigating data subject rights (e.g., the possibility for users to request the deletion of their personal data in mobile apps), and (iii) studying the impact of legal bases other than consent on developing compliant mobile apps.
@@ Acknowledgment
This research was funded in whole, or in part, by the Luxembourg National Research Fund (FNR), grant reference NCER22/IS/16570468/NCER-FT. For the purpose of open access, and in fulfillment of the obligations arising from the grant agreement, the author has applied a Creative Commons Attribution 4.0 International (CC BY 4.0) license to any Author Accepted Manuscript version arising from this submission.
[…]

### v2

@@ 5 Findings
  • addressing privacy considerations entailed from using third-party apps or libraries [28, 47],
  • evaluating the consistency between privacy policies and actual app code [37, 52, 54, 60],
  • detecting abuse of data collection due to granted permissions [42, 43, 49, 50]
@@
The aforementioned technical contributions were in all cases accompanied with validation elements such as, e.g., proof of concept examples [6], user studies [16, 59], or empirical validation using datasets. These datasets vary in size from small (with less than 50 elements such as apps, privacy policies, app reviews, or programming methods [42]), medium (51–500 elements [16, 28, 30, 50, 52, 53, 57]), to large (with more than 500 elements [1, 8, 20, 35, 36, 37, 41, 44, 46, 49, 51, 53, 54].)
As part of this RQ, we also looked into the app stores and operating systems targeted in the literature. Our results show that 32 papers considered Android only, seven considered both Android and iOS, one paper considered Android and Windows systems, and one more paper focused exclusively on iOS. The literature aligns with Android being the largest share of the mobile market compared to iOS. Studies covering both operating systems often rely on user surveys [11, 19, 9], governance or policy analyses [10, 18, 38], third-party consent management platforms [49], or screenshots of user interfaces [57], rather than focusing on concrete app code analyses tasks in relation to GDPR-related privacy concerns.
@@
By examining the 25 studies with technical contributions, we observe that all but four are exclusively focused on Android. Regarding the outliers, one utilizes Amazon Alexa (a virtual assistant) [37], one is based on iOS [59], and the last two mention Apple App Store as well as Google Play Store for work based on explicit consent (through GUI testing) and privacy policies [49, 57]. The remaining 21 studies focusing on Android include: One study that investigates Xiaomi market Store [43], five lack comprehensive details about app stores [41, 42, 44, 50, 60], and the remaining ones utilize GooglePlay [8, 16, 20, 28, 30, 35, 36, 41, 44, 46, 47, 51, 52, 53, 54].
In response to RQ3, we found that research contributions can be broadly categorized into two types: technical contributions, focusing on providing (semi-) automated support to address specific privacy concerns, and non-technical contributions, which are mostly based on user surveys. Most existing research focuses on the Android market and, to a limited extent, iOS. We also observed that several privacy concerns remain under-explored in the literature, calling for further research.
@@
Out of the 60 papers, only 14 papers ($\approx$23%) have shared some artifacts associated with the conducted research; Table 8 lists these papers and indicates the type of artifacts (source code, datasets, evaluation data files) made available. Out of these 14 papers, only two papers [37, 50] released all three types of artifacts.
Table 9 further investigates the licensing schema and on which type of platform is located the available material. We note that two studies [37, 46] have different licensing schemes regarding their source code and the additional material, which are stored separately in a different repository.
Regarding licenses, 5 of the 14 papers do not mention any explicit license and an additional 6th consider a vague license (other - open) leading to $\approx$43% of the work not being explicit about their licensing schemes. Regarding repositories, 11 of the 14 papers share (part of) their material through project webpages on personal website or GitHub repositories; only five consider persistent third-party repositories, including one study on the publisher’s page for additional content and four studies on Zenodo (with two papers sharing material in both Github and Zenodo).
@@
Source Code Datasets Results Papers
[37, 50]
[28, 46, 47]
[2]
[1, 43, 48, 53, 60]
[8, 21]
[3]
Table 9: Licensing and Sharing Approaches over Available Material
License type Total Papers
No license mentioned 5 [2, 21, 28, 48, 50]
Apache 2.0 1 [37]
CC0-1.0 license 1 [46]
Creative Commons Attribution 4.0 International 4 [3, 46, 47, 53]
CC BY 4.0 1 [8]
GNU GPL 3.0 2 [1, 60]
MIT License 1 [43]
other (open) 1 [37]
Sharing type Total Papers
Personal website 1 [28]
GitHub Project 10 [1, 2, 21, 37, 43, 46, 48, 50, 53, 60]
Publisher Repository 1 [3]
Open Science Repository 4 [8, 37, 46, 47]
@@ 5.5 Research Gaps
G1. More research is needed to better understand personal data collection practices in mobile apps. Table 6 shows that the primary studies have extensively investigated the collection of personal data and recipients of personal data sharing activities. These topics are often related to analyzing security issues concerning data leaks. To this end, we identify the following research gaps:
  • G1.1
@@
  • G1.2
    Nowadays, personal data collection typically involves permission requests to access personal data collected by other apps on the same mobile device. Another research gap concerns the investigation of indirect collection of personal data. It is important to understand what personal data is collected indirectly (e.g., scraping the web for data available online sources or accessing datasets through data brokers), how the process is documented, and whether users are aware of that.
  • G1.3
    The literature has carefully investigated personal data categories that are directly collected from individuals based on, e.g., consent. Personal data that describe users or their mobile devices, which are typically collected indirectly through other means are not very well explored in the literature. A key research gap is to investigate to what extent such personal data categories contain personal identifiable information and would therefore require special attention concerning GDPR compliance practices. For instance, the mobile IP address could reveal the geographical location of an individual. .
G2. Further research is needed to investigate legal bases beyond “consent” and their impact on the development of legally-compliant mobile apps. Our results show that most primary studies investigate the explicit consent of a user for data collection and processing. While consent is indeed an important legal basis, other legal bases are as important. For instance, collecting and processing personal data based on legitimate interest for the purpose of investigating a security attack is a good example where consent might not even be asked. The legal basis for handling personal data provides a complex, yet unexplored research landscape.
G3. There is a need for further research on the elicitation of comprehensive, implementable requirements pertaining to data subject rights. Our SLR highlights that the existing work investigated data subject rights and their implementation in mobile apps only to a very limited extent. According to the GDPR, users (as data subjects) should be able to easily exercise their rights to access their personal data, lodge a complaint, etc. For instance, mobile apps should provide adequate and explicit information (ideally through direct interactions) for users to request account deletion or to access all their personal data collected during a specific time span [50]. In-depth understanding of data subject rights should first be established in research to then make its way into practice. Without such an awareness level on the different users’ rights, developing better and legally-compliant apps is subject to the random, potentially correct interpretations of developers.
@@ 6 Threats to Validity
Internal Validity. The main consideration concerning the internal validity is selection bias. To mitigate the effect of this threat, we followed the guidelines reported in the literature for conducting systematic literature reviews. The results of our data extraction are also made publicly available and are thus open to scrutiny. The fact that we did not have a formal quality assessment checklist is another potential internal threat to validity, as it might lead to potential unequal treatment of primary studies. To mitigate this threat, we applied strict inclusion criteria based on the topic relevance, and we held several meetings among the authors to ensure shared understanding and minimize subjective judgment.
@@
External Validity. The findings of this study are primarily pertinent to GDPR-relevant privacy concerns in mobile app research, which may limit their generalizability to other regulations or software systems. We note that several identified concepts, such as informed consent, and corresponding research gaps, e.g., the need to explore legal bases beyond consent, may also be relevant across other research areas. However, we acknowledge that we provide no empirical evidence on their applicability beyond the scope of our study. To partially mitigate this threat, we explicitly defined the time range of interest for our SLR on the GDPR, iteratively refined our search query, and retrieved the primary studies from multiple online repositories and search engines. We further interpreted the results with caution and grounded our findings in existing research. We leave it to the research community to examine whether similar patterns can be identified in broader research contexts.
@@ 7 Conclusion
In this paper we have presented our systematic literature review of the research landscape in software engineering (SE) on addressing GDPR-relevant privacy concerns in mobile apps. We have reviewed 60 primary studies and analyzed them with respect to an existing conceptual model that provides a comprehensive set of concepts representing the GDPR.
We found that existing studies predominantly address three key GDPR-related privacy concepts: (i) the collection of personal data directly from users through mechanisms like cookies PD ORIGIN, (ii) the sharing of personal data with external entities (e.g., third parties) beyond the mobile apps (RECIPIENTS), and (iii) the acquisition of consent from users (LEGAL BASIS - CONSENT). Additionally, our results show that the majority of existing work investigates problems such as data leakage, app permission requests, and obtaining users’ consent. To address these challenges, existing work relies to a large extent on natural language processing, machine learning, and program analysis.
We further identified research gaps that highlight the need for further research in: (i) personal data collection, involving both checking the consistency between personal data categories disclosed in the app’s privacy policy against what the app actually collects as well as personal data collected indirectly, (ii) investigating the impact of legal bases beyond consent on the development of compliant mobile apps, and (iii) eliciting detailed implementable requirements pertinent to data subject rights.
@@ Acknowledgment
This research was funded in whole, or in part, by the Luxembourg National Research Fund (FNR), grant reference NCER22/IS/16570468/NCER-FT.